# meds_reader: A fast and efficient EHR processing library

### v1

@@ Abstract
This paper uses two electronic health record datasets, MIMIC-III and MIMIC-IV (Johnson et al., 2016, 2023), which are publicly available de-identified hospital, ICU, and ED data from Beth Israel Deaconess Medical Center. The code for this paper can be found at https://github.com/som-shahlab/meds_reader_paper_code.
@@
This research uses de-identified data and thus did not require IRB approval.
@@ 3 Design & Implementation
meds_reader is designed around the idea that electronic health record data is event stream data that consists of a collection of subjects, where each subject contains a sequence of time-stamped event data. EHR processing pipelines can then be written as a series of transformation steps of those per-subject event streams. Figure A shows the code for one example transformation that identifies all patients who have experienced a stroke. meds_reader uses a variety of optimizations to make these transformations as fast and memory-efficient as possible.
Python # Define the transformation def find_stroke(subjects): subject_ids = set() for subject in subjects: for event in subject.events: # For simplicity, only use ICD10/I63.* to define the phenotype if event.code.startswith(’ICD10/I63’): subject_ids.add(subject.subject_id) return subject_ids
# Apply the transformation over a database, using at most 10 threads subject_database = meds_reader.SubjectDatabase(’…’, num_threads=10) all_subjects = set() for partial_result in subject_database.map(find_stroke): all_subjects |= partial_result
@@
Python API: meds_reader is implemented as three main Python classes: Event, which represents individual measurements in EHR data; Subject, which contains all events for a single subject; and SubjectDatabase, which manages collections of subjects for efficient querying and processing. A complete overview of the API for this package can be found in Appendix C.
@@ 4 Benchmarks
We focus on reimplementing labeling and featurization for predicting the length of stay on both PyHealth (Yang et al., 2023) and MIMIC-IV Data Pipeline (Gupta et al., 2022). PyHealth and MIMIC-IV Data Pipeline were chosen as they are popular projects that use very different implementation strategies. PyHealth focuses more on Python pickle and CPU efficiency, while MIMIC-IV Data Pipeline relies more heavily on pandas and is more memory efficient. Labeling and featurization were chosen as reimplementation targets as they have concrete input/output expectations that are easy to verify. We perform experiments on the public datasets MIMIC-III and MIMIC-IV.
One complication with reimplementation experiments is that PyHealth and MIMIC-IV Data Pipeline use different subsets of MIMIC, including different tables and columns within those tables. In order to perform fair comparisons, we convert data from PyHealth and MIMIC-IV Data Pipeline’s internal formats to MEDS to make sure that meds_reader processes the same data as the original pipeline. This conversion would not have been necessary if PyHealth and MIMIC-IV Data Pipeline had been designed with MEDS support from the start. This conversion is primarily limited by the efficiency of PyHealth and MIMIC-IV Data Pipeline, but is relatively fast, with performance reported in Appendix A. We also need to convert from MEDS to meds_reader’s SubjectDatabase. Converting to a SubjectDatabase is a once-per-dataset cost that is reported in Appendix B.
We then reimplement labeling and featurization independently for both pipelines, and run the original code and our reimplementation against MIMIC-IV. Unfortunately, we cannot successfully run PyHealth on MIMIC-IV due to out-of-memory issues, so we instead use MIMIC-III for PyHealth experiments. We use GNU Time (GNU, ) to measure the wall clock time and peak memory usage. We measure disk usage by inspecting the cache directories for the pipelines. Appendix D contains further details and references to the code for these experiments. Table 1 contains the time, memory, and disk usage statistics for these experiments.
@@ 5 Discussion And Conclusion
However, there are some important limitations of meds_reader. First, meds_reader currently supports per-event properties that correspond directly to a single Python type. For example, we support strings, integers, floating point numbers, and datetimes, but do not support lists and structs. Events with list or struct properties require additional handling when using meds_reader, hurting readability and performance. We expect to address this in future versions by adding direct support for those types. The second limitation of meds_reader is that it still relies heavily on Python, as the user-supplied Python transformation functions are run directly in the Python interpreter. The reliance on Python, and the overhead caused by Python, require future work to convert the user-provided transformations from Python to native code to avoid that overhead.
@@ Appendix A Conversion of PyHealth and MIMIC-IV Data Pipeline to MEDS
| MIMIC-III | PyHealth | 0.87 | 4.45 |
Appendix C Package Documentation
See pages - of meds_reader.pdf
Appendix D Reimplementation Code And Setup

### v2

@@ Abstract
We use three publicly available electronic health record datasets: eICU, MIMIC-III and MIMIC-IV (Pollard et al., 2018; Johnson et al., 2016, 2023). Our paper code can be found at https://github.com/som-shahlab/meds_reader_paper_code.
@@
We use de-identified data and thus did not require IRB approval.
@@ 2 Related Work
Python def find_stroke(subjects): subject_ids = set() for subject in subjects: for event in subject.events: if event.code.startswith(’ICD10/I63’): subject_ids.add(subject.subject_id) return subject_ids
subject_database = meds_reader.SubjectDatabase(’…’, num_threads=10) all_subjects = set() for partial_result in subject_database.map(find_stroke): all_subjects |= partial_result
@@ 3 Design & Implementation
meds_reader is designed around the idea that electronic health record data is event stream data that consists of a collection of subjects, where each subject contains a sequence of time-stamped event data. EHR processing pipelines can then be written as a series of transformation steps of those per-subject event streams. Figure 2 shows the code for one example transformation that identifies all patients who have experienced a stroke. meds_reader uses a variety of optimizations to make these transformations as fast and memory-efficient as possible.
@@
Python API And Usage: meds_reader is implemented as three main Python classes: Event, which represents individual measurements in EHR data; Subject, which contains all events for a single subject; and SubjectDatabase, which manages collections of subjects for efficient querying and processing. A complete overview of the API for this package can be found at https://meds-reader.readthedocs.io/en/latest/api_reference.html. In order to use this API, it is necessary to first convert your dataset to MEDS, and then to the meds_readerSubjectDatabase, before finally using the resulting database for labeling and featurization (see Figure 3).
@@ 4 Benchmarks
We focus on reimplementing labeling and featurization for predicting the length of stay on both PyHealth (Yang et al., 2023) and MIMIC-IV Data Pipeline (Gupta et al., 2022). PyHealth and MIMIC-IV Data Pipeline were chosen as they are popular projects that use very different implementation strategies. PyHealth focuses more on Python pickle and CPU efficiency, while MIMIC-IV Data Pipeline relies more heavily on pandas and is more memory efficient. Labeling and featurization were chosen as reimplementation targets as they have concrete input/output expectations that are easy to verify. We perform experiments on the public datasets eICU, MIMIC-III and MIMIC-IV.
One complication with reimplementation experiments is that PyHealth and MIMIC-IV Data Pipeline use different subsets of MIMIC, including different tables and columns within those tables. In order to perform fair comparisons, we convert data from PyHealth and MIMIC-IV Data Pipeline’s internal formats to MEDS to make sure that meds_reader processes the same data as the original pipeline. This conversion would not have been necessary if PyHealth and MIMIC-IV Data Pipeline had been designed with MEDS support from the start. This conversion is primarily limited by the efficiency of PyHealth and MIMIC-IV Data Pipeline, but is relatively fast, with performance reported in Appendix A. We also need to convert from MEDS to meds_reader’s SubjectDatabase. Converting to a SubjectDatabase is a once-per-dataset cost that is reported in Appendix B. This once-per-dataset cost should not be important for repeated analysis on the same dataset, as it will be amortized across many runs, but might be important when trying to run one analysis across many datasets.
We then reimplement labeling and featurization independently for both pipelines, and run the original code and our reimplementation against MIMIC-IV. Unfortunately, we cannot successfully run PyHealth on MIMIC-IV due to out-of-memory issues, so we instead use MIMIC-III and eICU for PyHealth experiments. We use GNU Time (GNU, ) to measure the wall clock time and peak memory usage. We measure disk usage by inspecting the cache directories for the pipelines. Appendix C contains further details and references to the code for these experiments. Table 1 contains the time, memory, and disk usage statistics for these experiments.
@@ 5 Discussion And Conclusion
However, there are some important limitations of meds_reader. First, meds_reader currently supports per-event properties that correspond directly to a single Python type. For example, we support strings, integers, floating point numbers, and datetimes, but do not support lists and structs. Events with list or struct properties require additional handling when using meds_reader, hurting readability and performance. We expect to address this in future versions by adding direct support for those types. The second limitation of meds_reader is that it still relies heavily on Python, as the user-supplied Python transformation functions are run directly in the Python interpreter. The reliance on Python, and the overhead caused by Python, require future work to convert the user-provided transformations from Python to native code to avoid that overhead. A final limitation of meds_reader is that it depends on the MEDS standard and does not support data elements such as high-time resolution data (such as heart-rate monitors) that are not currently supported by MEDS. However, we expect this issue to be solved as more and more data support is added to MEDS.
@@ Appendix A Conversion of PyHealth and MIMIC-IV Data Pipeline to MEDS
Appendix C Reimplementation Code And Setup